sue can shovel snow from her driveway in 45 minutes. jum can do the same job in 40 minutes. how long would it take sue and jum to shovel the driveway if they worked together?
Answer by jim_thompson5910(35256)

1/45 + 1/40 = 1/t

360t*(1/45) + 360t*(1/40) = 360t*(1/t)

8t + 9t = 360

17t = 360

t = 360/17

t = 21.176

So if they worked together, it would take them approximately 21.176 minutes.